Solve the following equations.  1.6 = y/15 
Solve the following equations.  1.6 = y/15 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1.6 = y/1.5
y/1.5 = 1.6
y = 1.6 × 1.5
y=2.4

Thanks